# VICON: Vision In-Context Operator Networks for Multi-Physics Fluid Dynamics Prediction

### v1

@@ 1 Introduction
The In-Context Operator Network (ICON) is able to implicitly incorporate PDE system dynamics through input-output function pairs (before-after operators), achieving few-shot generalization without retraining or knowledge of the PDE information. As shown in (Yang et al., 2023a), a single ICON model can handle forward and inverse problems in various ODEs, PDEs, and mean-field control scenarios. However, vanilla ICONs face computational efficiency challenges: representing functions as scattered point tokens results in quadratic computational complexity relative to the number of data points. This scalability limitation has confined previous ICON applications to simple 1D problems, with only one instance of a 2D case using sparse data points. In this work, we propose a novel approach that leverages vision transformers for in-context operator learning of 2D functions. Our method represents 2D functions as grid-based images and divides them into sequences of patches. To enable “next function prediction” like in the original ICONs, we extend beyond classic vision transformers, where they process the sequence of patches of a single image, to handle sequences of input-output function pairs.
We name our new method Vision In-Context Operator Networks (VICON). The model architecture is illustrated in Figure 1. Our main contributions include the following.
@@
  • We evaluated our model’s performance in forward prediction tasks for different kinds of fluid dynamics system with varying timestep sizes, comprising 879K frames in total with diverse dynamics. Empirical results show better long-term prediction performance and lower computational cost compared to the state-of-the-art competitor: multi-physics pretraining (MPP) (McCabe et al., 2023).
  • VICON demonstrates advantages in the flexibility in forming function pairs, compared to strict sequence-to-sequence models, and the ability to adopt flexible inference strategies, such as multistep predictions.
@@ 3 Preliminaries
Rollout Rescaled $L^{2}$ Error Case Ours (single step) Ours (multi step) MPP-B (McCabe et al., 2023)
@@ 4.2 Vision In-Context Operator Networks
First, the input ${\bm{u}}_{t}$ and output ${\bm{u}}_{t+\Delta t}$ functions are divided into patches $\{{\bm{p}}_{t}\in\mathbb{R}^{R_{x}\times R_{y}\times c}\}_{k=1}^{N_{c}}$ and $\{{\bm{p}}_{t+\Delta t}\in\mathbb{R}^{R_{x}\times R_{y}\times c}\}_{l=1}^{N_{q}}$, where $R_{x},R_{y}$ are the resolution of a patch and $N_{c}$ and $N_{q}$ are the number of patches for the input and output functions, respectively. Note that $N_{q}$ can differ from $N_{c}$; for example, additional boundary paddings in the input functions result in $N_{c}>N_{q}$. However, in this work, we maintain $N_{c}=N_{q}$ throughout our experiments. For conciseness, we denote the $k$-th patch of the input function as ${\bm{C}}_{i}^{k}$ and the $l$-th patch of the output function as ${\bm{Q}}_{i}^{l}$, where $i\in\{1,\ldots,I\}$ is the pair index. These patches are then projected into a unified $d$-dimensional latent embedding space using a shared learnable linear function $f_{\phi}:R_{x}\times R_{y}\times c\rightarrow\mathbb{R}^{d}$:
@@
To support autoregressive prediction similar to Eq 2, VICON employs an alternating-sized (in the case when $N_{c}\neq N_{q}$) block-causal attention mask, as opposed to the conventional triangular causal attention mask as in mainstream generative large language models (Brown et al., 2020) and large vision models (Bai et al., 2024). The mask is defined as follows:
@@
After obtaining the output tokens from Eq 2, we extract the tokens corresponding to the input patches $\mathbf{c}_{i}^{1},\ldots,\mathbf{c}_{i}^{N_{c}}$ and denote them as $\tilde{{\bm{q}}}_{i}^{1},\ldots,\tilde{{\bm{q}}}_{i}^{N_{c}}$. These tokens are then projected back to the original physical space using a shared learnable linear function $g_{\psi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{R_{x}\times R_{y}\times c}$:
@@ 4.4 Datasets and Data Augmentation
The dataset used in this work contains three types of equation modeling different fluid dynamics:
@@ 4.5 Inference with Flexible Strategies
As mentioned in Section 4.4, VICON can make predictions with different timestep strides. Specifically, given $I_{0}$ initial frames $\{\mathbf{u}_{i}\}_{i=1}^{I_{0}}$, we can form in-context example pairs$\{\langle\mathbf{u}_{i},\mathbf{u}_{{i+s}}\rangle\}_{i=1}^{I_{0}-s}$. Then for $j\geq 1$, we can set the question condition (${\bm{c}}_{k}$ in Equation 3) as ${\bm{u}}_{j}$ to predict ${\bm{u}}_{j+s}$, i.e. $s$-step prediction.
To make predictions for long-term rollout, we can use VICON in an autoregressive fashion. The flexibility in predicting with different strides enables various rollout strategies. We explore two natural approaches: single-step rollout and multi-step rollout that advances in large strides to reduce the number of rollout steps.
For single-step rollout, we fix the in-context examples as $\{\langle\mathbf{u}_{i},\mathbf{u}_{{i+1}}\rangle\}_{i=1}^{I_{0}-1}$, and call the model to make prediction of $\mathbf{u}_{I_{0}+1}$ with $\mathbf{u}_{I_{0}}$ as the question condition. Then we can predict $\mathbf{u}_{I_{0}+2}$ with the predicted $\mathbf{u}_{I_{0}+1}$ as the question condition. So on and so forth. Multi-step rollout is more complicated: suppose the maximum prediction stride is $s_{\text{max}}$, we first make $s$-step predictions with $\mathbf{u}_{I_{0}}$ as the question condition, to make prediction of $\left\{\mathbf{u}_{I_{0}+s}\right\}_{s=1}^{s_{\text{max}}}$. Then with the predicted $\left\{\mathbf{u}_{I_{0}+s}\right\}_{s=1}^{s_{\text{max}}}$ as the question conditions, we make $s_{\text{max}}$-step predictions to get $\left\{\mathbf{u}_{I_{0}+s_{\text{max}}+s}\right\}_{s=1}^{s_{\text{max}}}$. So on and so forth. Such multi-step rollout strategy is the same as the one in (Yang & Osher, 2024). Examples of rollout strategies are provided in the Appendix B.3.
@@ 5 Experimental Results
The detailed experimental results are presented in Table 1 and Table 8, where we report the scaled and absolute $L^{2}$ roll-out errors for the three datasets. We record the rollout error at time steps 1, 5, 10, and the last step, along with the average error across all time steps. Due to the flexibility of time step strides, a single trained model adopts two rollout strategies: single-step and multi-step (Section 4.5) during inference, and we report results for both strategies.
Superior Accuracy for Long-term Predictions
We compare our model with MPP (McCabe et al., 2023), which is trained as a single-step predictor that predicts the next frame given the previous frames with a fixed $\Delta t$. We observe that VICON (both strategies) outperforms MPP at all timesteps for the PDEBench-Comp-LowVis and PDEBench-Comp-HighVis datasets. For longer rollouts, VICON shows a significant reduction in scaled $L^{2}$ error of $40\%$ and $61.6\%$ for PDEBench-Comp-LowVis and PDEBench-Comp-HighVis datasets, respectively. For the PDEArena-Incomp dataset, while VICON slightly underperforms MPP initially, it achieves better results for longer-step rollouts (the last step and the averaged error across the whole time horizon of 46 steps). In general, our method demonstrates superior long-term predictions compared to MPP.
Flexible Rollout Strategy
@@
This performance difference can be attributed to two key factors: (1) PDEBench-Comp-LowVis and PDEBench-Comp-HighVis employ significantly larger recording timestep sizes with fewer total frames (detailed in Section A), making operator learning more challenging. In such cases, the error in a single large step outweighs the potential benefits of reduced autoregressive rollout numbers. (2) Conversely, PDEArena-Incomp contains longer trajectories with smaller recording timesteps, making operator learning easier, where larger strides reduce the number of rollouts and thus error accumulation.
@@
Stride Generalization
@@
Impact of Data Diversity on Generalization
@@
Resource and Timing Metrics Ours MPP-B (McCabe et al., 2023)
@@
(a) 5 output steps for PDEBench-Comp-HighVis dataset. The channel plotted is the pressure field in equation 12. Each column represents a different timestamp.
(b) 5 output steps for PDEArena-Incompdataset. The channel plotted is the x-velocity in equation 9, i.e. $u_{x}$. Each column represents a different timestamp.
Figure 3: Two example outputs for the VICON model.
@@ 6 Conclusion and Future Works
Several promising directions remain for future investigation. Our method currently requires full domain input. 1. Incorporating physical priors, such as translational invariance, could enable domain decomposition for training on smaller subdomains, effectively decoupling computational complexity from domain size. 2. Extending VICON to handle irregular domains and different data structures such as graphs or point clouds would broaden its applicability to more realistic scenarios. 3. Our current approach of joining channels across datasets has limited scalability. Drawing inspiration from machine translation, where subword tokenization enables handling multiple languages without a fixed vocabulary union, developing similar adaptive approaches for physical field representations could offer a more scalable solution.
@@ B.1 VICON Model Details
Input patch dimension $8\times 8$
Output patch dimension $8\times 8$
@@ B.3 Rollout Strategy Example
For a fair comparison, we retrained MPP-B using our training dataset using the same model configurations and optimizer hyperparameters. We evaluate MPP-B using the same testing setup and metric.
@@ Appendix C More Results and Visualizations
Table 8 summarizes the rollout $L^{2}$ errors, and Figure 4 presents additional visualizations of the VICON model outputs.
@@
(a) The channel plotted is the y-velocity.
(b) The channel plotted is the density field.
(c) The channel plotted is the pressure field.
Figure 4: Example outputs for the VICON model. 5 output steps for PDEBench-Comp-LowVis dataset. Each column represents a different timestep.

### v2

# VICON: A Foundation Model for Multi-Physics Fluid Dynamics via Vision In-Context Operator Networks

@@ 1 Introduction
While these approaches require explicit encoding of PDE information, which may not be readily available or precisely known in practice, the In-Context Operator Network (ICON) (Yang et al., 2023a) offers a fundamentally different approach. ICON implicitly encodes system dynamics through input-output function pairs (before-after operators), achieving a few-shot generalization without retraining or explicit PDE knowledge. A single ICON model has shown success in handling both forward and inverse problems in various ODEs, PDEs, and mean field control scenarios.
However, ICON’s computational efficiency becomes a critical bottleneck when handling dense data in higher dimensions. Specifically, representing each spatial point as an individual token leads to quadratic complexity in the number of points, making it computationally prohibitive for practical 2D and 3D applications - to date, only one instance of a 2D case (Yang et al., 2023a) using sparse data points has been demonstrated.
To address this limitation, we propose Vision In-Context Operator Networks (VICON), which leverage vision transformers to process 2D functions using an efficient patch-wise approach. Going beyond classic vision transformers, which typically process patches from a single image, VICON extends the architecture to handle sequences of input-output function pairs, allowing “next function prediction” capabilities while preserving the benefits of in-context operator learning. Our contributions include:
@@
  • We evaluated our model across diverse fluid dynamics systems, comprising 879K frames with varying timestep sizes. Empirical results demonstrate significant improvements over the state-of-the-art MPP (McCabe et al., 2023), including a 40% and 61. 6% reduction in the scaled L2 error for long-term predictions in two benchmark datasets, a 67% reduction in the prediction error of turbulence kinetic energy, and approximately one third of the singe step inference time.
  • VICON enables flexible rollout scheme unattainable by traditional sequence-to-sequence models: it can form example pairs from nonsequential time sequences, handle scenarios with insufficient example frames, and adapt to varying timestep strides without retraining. These capabilities allow for robust rollout even with incomplete or irregularly sampled temporal data.
@@ 3 Preliminaries
Rollout Rescaled $L^{2}$ Error Case Ours (single step) Ours (multi step) MPP (McCabe et al., 2023)
@@ 4.2 Vision In-Context Operator Networks
First, the input ${\bm{u}}_{t}$ and output ${\bm{u}}_{t+\Delta t}$ functions are divided into patches $\{{\bm{C}}^{k}\in\mathbb{R}^{R_{x}\times R_{y}\times c}\}_{k=1}^{N_{c}}$ and $\{{\bm{Q}}^{l}\in\mathbb{R}^{R_{x}\times R_{y}\times c}\}_{l=1}^{N_{q}}$, where $R_{x},R_{y}$ are the resolution dimensions of the patch and $N_{c}$ and $N_{q}$ are the number of patches for the input and output functions, respectively. While $N_{q}$ can differ from $N_{c}$ (for instance, when input functions require additional boundary padding, resulting in $N_{c}>N_{q}$), we maintain $N_{c}=N_{q}$ throughout our experiments. For notational clarity, we add the subscript $i\in\{1,\ldots,I\}$ to denote the pair index, where ${\bm{C}}_{i}^{k}$ represents the $k$-th patch of the input function and ${\bm{Q}}_{i}^{l}$ represents the $l$-th patch of the output function in the $i$-th pair. These patches are then projected into a unified $d$-dimensional latent embedding space using a shared learnable linear function $f_{\phi}:R_{x}\times R_{y}\times c\rightarrow\mathbb{R}^{d}$:
@@
To support autoregressive prediction similar to Eq 2, VICON employs an alternating-sized (in the case where $N_{c}\neq N_{q}$) block causal attention mask, as opposed to the conventional triangular causal attention mask as in mainstream generative large language models (Brown et al., 2020) and large vision models (Bai et al., 2024). The mask is defined as follows:
@@
After obtaining the output tokens from Eq 2, we extract the tokens corresponding to the input patch indices $\mathbf{c}_{i}^{1},\ldots,\mathbf{c}_{i}^{N_{c}}$ and denote them as $\tilde{{\bm{q}}}_{i}^{1},\ldots,\tilde{{\bm{q}}}_{i}^{N_{c}}$. These tokens are then projected back to the original physical space using a shared learnable linear function $g_{\psi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{R_{x}\times R_{y}\times c}$:
@@ 4.4 Datasets and Data Augmentation
The dataset used in this work contains three types of equations modeling different fluid dynamics:
@@ 4.5 Inference with Flexible Strategies
As mentioned in Section 4.4, VICON can make predictions with varying timestep strides. Given $I_{0}$ initial frames $\{\mathbf{u}_{i}\}_{i=1}^{I_{0}}$, we can form in-context example pairs $\{\langle\mathbf{u}_{i},\mathbf{u}_{{i+s}}\rangle\}_{i=1}^{I_{0}-s}$. For $j\geq 1$, we set the question condition (${\bm{c}}_{k}$ in Equation 3) as ${\bm{u}}_{j}$ to predict ${\bm{u}}_{j+s}$, enabling $s$-step prediction.
For long-term rollout, we employ VICON in an autoregressive fashion. The ability to predict with different timestep strides enables various rollout strategies. We explore two natural approaches: single-step rollout and multi-step rollout, where the latter advances in larger strides to reduce the number of rollout steps.
For single-step rollout, we simply follow the autoregressive procedure with $s=1$. Multi-step rollout involves a more sophisticated approach: given a maximum prediction stride $s_{\text{max}}$, we first make sequential predictions with a gradually growing $s=1$ to $s_{\text{max}}$ using $\mathbf{u}_{I_{0}}$ as the question condition, obtaining $\left\{\mathbf{u}_{I_{0}+s}\right\}_{s=1}^{s_{\text{max}}}$. Using each frame in this sequence as a question condition, we then make consistent $s_{\text{max}}$-step predictions while preserving all intermediate frames. This multi-step strategy follows the approach in (Yang & Osher, 2024). Detailed examples of both rollout strategies are provided in Appendix B.3.
@@ 4.6 Evaluation Metric
(a) PDEBench-Comp-LowVis
(b) PDEBench-Comp-HighVis
(c) PDEArena-Incomp
Figure 3: Comparison of rollout errors using different patch resolutions ($8,16,32,64$).
@@ 5 Experimental Results
The detailed experimental results are presented in Table 1 and Table 8, where we report the scaled and absolute $L^{2}$ roll-out errors for the three datasets. We record the rollout error at time steps 1, 5, 10, and the last step, along with the average error across all time steps. Due to the flexibility of time steps, a single trained model adopts two rollout strategies: single-step and multistep (Section 4.5) during inference, and we report results for both strategies.
We benchmark against MPP (McCabe et al., 2023) (with the AViT-B model architecture, see details in their Table 3), the current state-of-the-art sequence-to-sequence model. The vanilla ICON, which would require processing over 114K tokens per frame (128×128×7), exceeds available GPU memory and is thus computationally infeasible for direct comparison on these datasets.
Superior Accuracy for Long-term Predictions.
We observe that VICON (both strategies) outperforms MPP at all timesteps for the PDEBench-Comp-LowVis and PDEBench-Comp-HighVis datasets. For longer rollouts, VICON shows a significant reduction in scaled $L^{2}$ error of $40\%$ and $61.6\%$ for PDEBench-Comp-LowVis and PDEBench-Comp-HighVis datasets, respectively. For the PDEArena-Incomp dataset, while VICON slightly underperforms MPP initially, it achieves better results for longer-step rollouts (the last step and the averaged error across the whole time horizon of 46 steps). In general, our method demonstrates superior long-term predictions compared to MPP.
Turbulence Kinetic Energy Analysis.
The turbulence kinetic energy (TKE), defined as $\frac{1}{2}(\overline{\tilde{u}_{x}^{2}}+\overline{\tilde{u}_{y}^{2}})$, is a critical metric to quantify the ability of a model to capture turbulent flow characteristics. Here, $\tilde{{\bm{u}}}={\bm{u}}-\overline{{\bm{u}}}$ denotes the fluctuation of the velocity from its statistical equilibrium state $\overline{{\bm{u}}}$. Within our datasets, only a subset of PDEBench-Comp-LowVis, specifically those initialized with fully developed turbulent fields (see Appendix D.5 of (Takamoto et al., 2022)), is suitable for TKE analysis. After filtering these entries, we compare the mean absolute error (MAE) of the TKE between VICON and MPP. Our approach achieves a TKE error of $0.016$, significantly outperforming the MPP’s error of $0.049$, which demonstrates superior preservation of turbulent flow characteristics. Figure 4 illustrates a representative comparison of TKE predictions between VICON and MPP.
Flexible Rollout Strategy.
@@
This performance difference stems from the datasets’ temporal characteristics: PDEBench-Comp-LowVis and PDEBench-Comp-HighVis use larger timesteps with fewer total frames, making accurate large-stride predictions challenging and favoring single-step approaches. In contrast, PDEArena-Incomp’s longer trajectories with smaller timesteps benefit from multi-step prediction, as reducing the number of rollout steps helps minimize error accumulation.
@@
Figure 4: Comparison of turbulence kinetic energy predictions between VICON and MPP, demonstrating the improved accuracy in capturing turbulent flow features of VICON. (Top-left) Ground truth TKE field, (Left) VICON prediction and error, (Right) MPP prediction and error.
Stride Generalization.
@@
Impact of Patch Resolutions.
We conducted ablation studies on patch resolution by varying patch sizes (8, 16, 32, 64). Although smaller patches could theoretically capture finer spatial structures, they generate longer token sequences for the transformer. Due to the quadratic attention complexity, this imposes a critical trade-off: we must reduce the token dimension to stay within memory constraints. Our experiments show that reducing the patch resolution below $16\times 16$ requires lowering both the token dimension from $1024$ to $512$ and the feedforward dimension from $2048$ to $1024$ to fit within our GPU memory limits. Moreover, further reduction below the $8\times 8$ patches proved infeasible even with mixed precision training.
The results in Figure 3 demonstrate that our choice of patch size 16 achieves optimal performance in all data sets, indicating an effective balance between the length of the sequence and the representational capacity. Although performance degradation with coarser patches (32, 64) aligns with theoretical expectations, the degradation with finer patch resolution probably stems from the necessary reduction in hidden dimensions. This observation highlights the importance of carefully balancing the sequence length against hidden dimensions. The substantial computational resources required also suggest a fundamental challenge in scaling VICON to 3D applications without stronger computing infrastructure or modifications to the model.
Impact of Data Diversity on Generalization.
@@
Resource and Timing Metrics Ours MPP (McCabe et al., 2023)
@@ 6 Conclusion and Future Works
Several promising directions remain for future investigation. Our method currently requires full domain input. 1. Incorporating physical priors, such as translational invariance, could enable domain decomposition for training on smaller subdomains, effectively decoupling computational complexity from domain size. 2. Extending VICON to handle irregular domains and different data structures such as graphs or point clouds would broaden its applicability to more realistic scenarios. 3. Our current approach of joining channels across datasets has limited scalability. Drawing inspiration from machine translation’s subword tokenization, developing adaptive physical field representations could offer a more scalable cross-dataset solution. 4. Most critically, extending to 3D applications faces fundamental challenges as token sequence length grows cubically with spatial dimensions, while memory constraints limit the hidden dimensions and representational capacity. Novel architectures combining in-context learning with efficient data compression techniques may be necessary to address this limitation.
@@ Acknowledgment
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ B.1 VICON Model Details
Input patch numbers $8\times 8$
Output patch numbers $8\times 8$
@@ B.3 Rollout Strategy Example
For a fair comparison, we retrained MPP using our training dataset using the same model configurations and optimizer hyperparameters. We evaluate MPP using the same testing setup and metric.
@@ Appendix C More Results and Visualizations
Table 8 summarizes the rollout $L^{2}$ errors; Figure 5 and Figure 6 present additional visualizations of the VICON model outputs.
@@
(a) A sequence of 5 output steps for PDEBench-Comp-HighVis dataset. The channel plotted is the pressure field in equation equation 12.
(b) A sequence of 5 output steps for PDEArena-Incomp dataset. The channel plotted is the x-velocity in equation equation 9.
Figure 5: Example outputs for the VICON model. (a) The pressure field of PDEBench-Comp-HighVis dataset, and (b) the x-velocity field of PDEArena-Incomp dataset. Each column represents a different timestep.
(a) The channel plotted is the y-velocity field in equation equation 9.
(b) The channel plotted is the density field in equation equation 9.
(c) The channel plotted is the pressure field in equation equation 9.
Figure 6: More example outputs for the VICON model. Showing 5 output steps for the PDEBench-Comp-LowVis dataset as governed by equation equation 9: (a) y-velocity, (b) density, and (c) pressure fields. Each column represents a different timestep.